OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of ? a. 5x – 3y = –19 b. 5x – 3y = 19 c. 3x – 5y = 19 d. 3x – 5y = –19

OpenStudy (anonymous):

***with a slope of 3/5?

OpenStudy (anonymous):

The right answer is c: 3x-5y=19 Write the equation first using the point and slope. y+2=(3/5)(x-3) Now expand: y+2=(3/5)x-(9/5). Now gather all y's and x's on one side of the equality. y-(3/5)x=-19/5. Multiply by 5 both sides to eliminate the fractions. 5y-3x=-19. Multiply both sides by -1 to get, -5y+3x=19. This is answer c, which you can write 3x-5y=19.